# Input Reduction Enhanced LLM-based Program Repair

### v1

@@ 6. Threats to Validity
The inherent stochasticity of LLMs may pose threats to internal validity. First, reducer generation is made deterministic by decoding with temperature $=0$; each prompt therefore yields identical code. Second, patch generation retains temperature $=0.8$ so that the model explores diverse fixes. Because developers typically review more than one suggestion, we sample k candidate patches per bug and evaluate pass@1, pass@5, and pass@10. Summarizing results through these statistics converts raw sampling noise into a controlled, reproducible measure, thereby mitigating the threat that randomness alone drives the observed accuracy.
@@
One concern is that ReduceFix might benefit only the pipeline evaluated in this study and fail to transfer to other APR frameworks. To investigate this limitation, we inserted the reducer as a plug‑in into ChatRepair while leaving the rest of its conversation logic unchanged. The modified system achieved a higher pass@10 on the same benchmark, indicating that the method can be integrated into existing APR tools with minimal engineering effort, mitigating the threat of limited applicability.
@@ 7.1. LLM-based Program Repair
Program Repair focuses on automatically or semi-automatically fixing software bugs. It aims to reduce the cost and effort of manual debugging by generating patches that correct faulty behavior in code. LLMs have recently advanced automated program repair, significantly surpassing traditional rule-based or search-based techniques (Yin et al., 2024; Yang et al., 2024; Xia and Zhang, 2022; Yang et al., 2025b, a). ChatRepair is the first work that leverages detailed feedback for each and every patch validated for conversational APR (Xia and Zhang, 2024). Following it, many LLM-based repair pipelines embed the failing test case in the prompt to supply bug context (Yang et al., 2024; Kong et al., 2024; Tang et al., 2024). However, when the failure-inducing test case is long, the LLMs’ context window is exceeded and attention diffuses, producing the “lost-in-the-middle” effect and lowering patch accuracy (Yang et al., 2024; Tian et al., 2023). Reducing the failure-inducing input while preserving the failure is therefore crucial for efficient, focused repair, especially when the goal is to fine-tune compact LLMs on highly informative examples. However, no existing work studies test input reduction in the context of APR pipelines. To our knowledge, our work ReduceFix is the first approach that leverages test input reduction techniques into LLM-based APR.
@@ 7.2. Test Input Reduction
Reducing test inputs that trigger bugs is crucial for efficient debugging. Delta debugging is the most popular approach for this purpose in software engineering. Zeller and Hildebrandt initially proposed the first delta debugging algorithm, named ddmin (Zeller, 1999; Zeller and Hildebrandt, 2002). Following ddmin, HDD (Misherghi and Su, 2006) and Perses (Sun et al., 2018) utilize the syntactical structure of the test input to further improve the reduction process. ProbDD (Wang et al., 2021) introduces a probabilistic model to improve ddmin, by estimating the probability of each element being kept in the produced result and prioritizing the reduction of those with high probabilities. GReduce (Ren et al., 2025) assumes that the given input is generated by a test generator and applies ddmin to the generator’s execution trace to reduce the test input. Besides the above delta debugging and its derived algorithms, many domain-specific approaches have been proposed for test input reduction on various domains. For example, CReduce (Regehr et al., 2012) reduces C/C++ programs by iteratively applying pre-defined, well-crafted program transformation rules. ddSMT (Niemetz and Biere, 2013) is proposed for conducting delta debugging on SMT formulas. Binary Reduction (Kalhauge and Palsberg, 2019) is proposed to reduce Java bytecode.
Performing reduction on various types of input formats, such as those in our benchmark tasks, requires manually customizing delta debugging algorithms or designing domain-specific approaches, which is time-consuming and requires domain expertise. Our work addresses this bottleneck by prompting an LLM to automatically generate the input reducer, enabling task-agnostic input minimization that feeds reduced tests directly back to the repair LLMs.
@@ 8. Conclusion
Our study addresses two persistent gaps that exist between test-case reduction and LLM-driven automated program repair. First, we demonstrate that LLMs can generate a task-specific reducer from a single prompt, thereby freeing developers from the need for manual customization to accommodate heterogeneous input formats. Second, we place this reduction step within the LLM-based repair loop so that the trimmed failing case provides a precise, high-signal prompt for patch generation. Experiments on the newly built LFTBench benchmark confirm the value of ReduceFix: inputs shrink by up to 100%, and repair success climbs by up to 53.8%. Ablation study shows that the benefit of ReduceFix stems from the combination of brevity and complete failure evidence rather than length trimming alone. An extension experiment further shows that replacing the test input in ChatRepair with the reduced one increases its pass@10 by 21.3%. These improvements demonstrate that integrating input reduction with LLM-based APR materially advances repair effectiveness.
Fully automated reduction unlocks practical applications: programming courses can return minimal counterexamples with fixes to accelerate students’ learning, and continuous-integration pipelines can store smaller regression tests and pinpoint faults for developers more efficiently. Because the reduced input is a plug‑in component, future work can extend ReduceFix into other APR frameworks and even broadly long‑context LLM tasks that profit from concise but information‑rich prompts.

### v2

@@ 6. Threats to Validity
The inherent stochasticity of LLMs may pose threats to internal validity. First, reducer generation is made deterministic by decoding with temperature =0=0; each prompt therefore yields identical code. Second, patch generation retains temperature =0.8=0.8 so that the model explores diverse fixes. Because developers typically review more than one suggestion, we sample k candidate patches per bug and evaluate pass@1, pass@5, and pass@10. Summarizing results through these statistics converts raw sampling noise into a controlled, reproducible measure, thereby mitigating the threat that randomness alone drives the observed accuracy.
@@
One concern is that ReduceFix might benefit only the pipeline evaluated in this study and fail to transfer to other APR frameworks. To investigate this limitation, we inserted the reducer as a plug-in into ChatRepair and CREF while leaving the rest of its conversation logic unchanged. The modified system achieved a higher pass@10 on the same benchmark, indicating that the method can be integrated into existing APR tools with minimal engineering effort, mitigating the threat of limited applicability. Another concern is about cross-language and scenario generalizability. We validated ReduceFix on LFTBench-Py and on repository-level OSS-Fuzz and observed improved repair performance in both settings.
@@ 7.1. LLM-based Program Repair
Program Repair focuses on automatically or semi-automatically fixing software bugs. It aims to reduce the cost and effort of manual debugging by generating patches that correct faulty behavior in code. LLMs have recently advanced automated program repair, significantly surpassing traditional rule-based or search-based techniques (yin_thinkrepair_2024; yang2024cref; xia2022less; yang2025enhancing; yang2025surveyllmbasedautomatedprogram; fan2023automated). ChatRepair is the first work that leverages detailed feedback for each and every patch validated for conversational APR (xia2024automated). Following it, many LLM-based repair pipelines embed the failing test case in the prompt to supply bug context (yang2024cref; kong2024contrastrepair; tang2024code). However, when the failure-inducing test case is long, the LLMs’ context window is exceeded and attention diffuses, producing the “lost-in-the-middle” effect and lowering patch accuracy (yang2024cref; tian2023chatgpt). Reducing the failure-inducing input while preserving the failure is therefore crucial for efficient, focused repair, especially when the goal is to fine-tune compact LLMs on highly informative examples. However, no existing work studies test input reduction in the context of APR pipelines. To our knowledge, our work ReduceFix is the first approach that leverages test input reduction techniques into LLM-based APR.
@@ 7.2. Test Input Reduction
Reducing test inputs that trigger bugs is crucial for efficient debugging. Delta debugging is the most popular approach for this purpose in software engineering. Zeller and Hildebrandt initially proposed the first delta debugging algorithm, named ddmin (ODD; DD). Following ddmin, HDD (HDD) and Perses (Perses) utilize the syntactical structure of the test input to further improve the reduction process. ProbDD (ProbDD) introduces a probabilistic model to improve ddmin, by estimating the probability of each element being kept in the produced result and prioritizing the reduction of those with high probabilities. GReduce (GReduce) assumes that the given input is generated by a test generator and applies ddmin to the generator’s execution trace to reduce the test input.
Besides the above delta debugging and its derived algorithms, many domain-specific approaches have been proposed for test input reduction on various domains. For example, CReduce (CReduce), Vulcan (10.1145/3586049), and T-Rec (10.1145/3690631) reduce source code by iteratively applying pre-defined, well-crafted program transformation rules. LPR (10.1145/3650212.3652126) and SimpT5 (wang2025towards) also focus on source code reduction, proposing an LLM-driven technique based on semantic-preserving program transformations. ddSMT (niemetz2013ddsmt) is proposed for conducting delta debugging on SMT formulas. Binary Reduction (DBLP:conf/sigsoft/KalhaugeP19) is proposed to reduce Java bytecode. Those approaches are designed for specialized scenarios in which the test inputs take specific forms, including source code and SMT formulas, and are thus not applicable to the competition problems and OSS-Fuzz projects used in our study.
To perform reduction on various types of input formats, such as those in our benchmark tasks, requires manually customizing delta debugging algorithms or designing domain-specific approaches, which is time-consuming and requires domain expertise. In contrast to domain-specific approaches, our work is applicable to a broader range of input types by leveraging LLM’s ability to understand specification descriptions. Our approach prompts an LLM to automatically generate the input reducer, enabling task-agnostic input minimization that feeds reduced tests directly back to the repair LLMs.
@@ 8. Conclusion
Our study addresses two persistent gaps that exist between test-case reduction and LLM-driven automated program repair. First, we demonstrate that LLMs can generate a task-specific reducer from a single prompt, thereby freeing developers from the need for manual customization to accommodate heterogeneous input formats. Second, we place this reduction step within the LLM-based repair loop so that the trimmed failing case provides a precise, high-signal prompt for patch generation. Experiments on the newly built LFTBench benchmark confirm the value of ReduceFix: inputs shrink by up to 100%, and repair success climbs by up to 53.8%. Ablation study shows that the benefit of ReduceFix stems from the combination of brevity and complete failure evidence rather than length trimming alone. An extension experiment further shows that replacing the test input in ChatRepair with the reduced one increases its pass@10 by 21.3%. These improvements demonstrate that integrating input reduction with LLM-based APR materially advances repair effectiveness. Evaluation on OSS-Fuzz projects further shows that ReduceFix can improve repair accuracy on industrial-scale projects.
Fully automated reduction unlocks practical applications: programming courses can return minimal counterexamples with fixes to accelerate students’ learning, and continuous-integration pipelines can store smaller regression tests and pinpoint faults for developers more efficiently. Because the reduced input is a plug-in component, future work can extend ReduceFix into other APR frameworks and even broadly long-context LLM tasks that profit from concise but information-rich prompts.
Acknowledgements.
This work has been partly supported by National Natural Science Foundation (Grant Number 62273292), China; by Central Leading Local Science and Technology Development Project of Hebei Province (Grant Number 246Z0804G), China; by Innovation Capability Improvement Plan Project of Hebei Province (22567626H), China.